OpenStudy (anonymous):

-24=4x-12 Does x= -3 ?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

That is correct

OpenStudy (anonymous):

of course x=-3 -24=4x-12 -24+12=4x -12=4x -3=x

OpenStudy (anonymous):

yups u got it right cheer up!!!!!!!!!!!